determine the magnitude of an electric field at r=0.01m from the center of non-conducting sphere with a radius R=0.001 possessing a charge of 7C which evenly distributed throughout the sphere. (r>R).


1
Expert's answer
2022-02-20T15:45:31-0500

"E=\frac{kqr}{R^3}=63\cdot 10^{16}~\frac NC."